OpenStudy (anonymous):

write an equation in slope-intercept form of the line whose y-intercept is -3 and the line is perpendicular line whose equation is 3x-2y=9 what is the slope intercept form? i need step by step help please

OpenStudy (anonymous):

slope intercept form means y=mx+n m is the slope n the y-intercept n=-3 you still need m you know that if 2 lines are perpendicular ,the product of their slopes is -1 first thing now is to determine the slope of 3x-2y=9

OpenStudy (anonymous):

i get -2y=-3/-2x-9/-2 correct

OpenStudy (anonymous):

then my final answer should be y=-3/-2x-3 correct

OpenStudy (anonymous):

if you have 2y=-3x+9 then divide all by 2 you get either y=(-3/2)x+(9/2) or you can factor to 3(x-3)/2 \[\frac{ 3(x-3)}{ 2}\] I prefer the first one which would look like \[-\frac{ 3 }{ 2 }x+\frac{ 9 }{ 2 }\]

OpenStudy (anonymous):

sorry I messed that up hang on a second

OpenStudy (anonymous):

oops 3/2x+9/2

OpenStudy (anonymous):

and okay

OpenStudy (anonymous):

yeah so it should be (-3x+9)/-2 which is \[\frac{ 3 }{ 2 }x-\frac{ 9 }{ 2 }\]

OpenStudy (anonymous):

okay and the final answer should be y=-3/2x-3 or y=-2/3x-3?

OpenStudy (anonymous):

@heradog ya there lol

OpenStudy (anonymous):

sorry trying to do one of mine. yes y=-2/3x-3

OpenStudy (anonymous):

because your perpendicular line is a negative reciprocal of the original one so the original was (3/2) and the negative reciprocal is (-2/3)

OpenStudy (anonymous):

and they give you the y intercept of -3 or b

OpenStudy (anonymous):

thanks herdog

OpenStudy (anonymous):

no problem